# Unraveling the Spin Coulomb Drag Effect and Its Impact on Spin Transport in the Three-Dimensional Electron Gas

### v1

@@ Abstract
The spin Coulomb drag effect, arising from the exchange of momentum between electrons of opposite spins, plays a crucial role in the spin transport of interacting electron systems. This effect leads to the emergence of a spin mass and the finite lifetime of spin currents, posing challenges for the accurate description of spin dynamics. Using the state-of-the-art Variational Diagrammatic Monte Carlo approach, we investigate the spin-resolved exchange-correlation (XC) kernel in the three-dimensional uniform electron gas. Our analysis reveals a distinct nature in the spin response, characterized by a $1/q^{2}$ divergence in the spin XC kernel at finite frequencies. This so-called ultranonlocal behavior, stemming from the spin Coulomb drag effect, is absent in the charge channel. By extracting precise values for the spin mass enhancement factor, we observe a trend of increasing enhancement with decreasing electron density. Furthermore, we find compelling evidence for the suppression of spin diffusion at low temperatures, characterized by vanishing inverse relaxation times. These findings deepen our understanding of the intricate relation between the Coulomb interaction and the spin transport, providing valuable insights for the development of accurate density functional approximations and the advancement of spintronics and quantum technologies.
@@ I Introduction
Spin transport in interacting electron systems is a fundamental problem with far-reaching implications for spintronics and quantum technologies. The key challenge in describing spin transport lies in the complex interplay between electron-electron interactions and the non-conservation of spin currents. Unlike charge currents, which are conserved, spin currents can be dissipated through the exchange of momentum between electrons of opposite spins. This dissipation mechanism, known as the spin Coulomb drag effect D’Amico and Vignale (2000); D’Amico and Ullrich (2010), leads to a friction between the spin currents of different components and gives rise to two major consequences: the generation of a spin mass and the finite lifetime of spin currents.
The spin Coulomb drag effect manifests itself through the emergence of a spin mass $m_{s}$, a many-body parameter that determines the spin current carried by a single quasiparticle Qian et al. (2004). In the presence of electron-electron interactions, the spin current carried by a quasiparticle of momentum $\mathbf{p}$ and spin $\sigma$ is proportional to $\mathbf{p}/m_{s}$, rather than $\mathbf{p}/m$, where $m$ is the bare electron mass. The spin mass $m_{s}$ is larger than both the bare electron mass and the quasiparticle effective mass $m^{*}$, indicating that interactions have a significant impact on spin transport.
In addition to the spin mass, the spin Coulomb drag effect leads to a dissipation of spin currents, namely the spin diffusion phenomenon, characterized by a relaxation time $\tau_{sd}$ D’Amico and Vignale (2000). This relaxation time quantifies the lifetime of spin currents in the presence of electron-electron interactions and plays a crucial role in determining the spin diffusion length and the efficiency of spin transport.
The spin Coulomb drag effect and its consequences have profound implications for the theoretical description of spin transport. In particular, they give rise to a singular structure in the spin-resolved exchange-correlation (XC) kernel of the spin response function Qian et al. (2003); Qian and Vignale (2003). The spin XC kernel denoted as $K_{\mathrm{\rm XC}}^{-}(\mathbf{q},\omega)$, is defined through the spin-spin response function $\chi_{S}(\mathbf{q},\omega)$ and the non-interacting response function $\chi_{0}(\mathbf{q},\omega)$ via the Dyson equation Giuliani and Vignale (2005):
$\chi_{S}(\mathbf{q},\omega)=\frac{\chi_{0}(\mathbf{q},\omega)}{1-K_{\mathrm{\rm XC}}^{-}(\mathbf{q},\omega)\chi_{0}(\mathbf{q},\omega)}$ (1)
Here, the superscript ‘--’ in $K_{\mathrm{\rm XC}}^{-}$ denotes the antisymmetric XC kernel related to the spin channel, while the symmetric XC kernel $K_{\rm XC}^{+}$ related to the charge channel will be discussed in the Eq. (5). The spin XC kernel encodes the effects of electron-electron interactions on the spin-spin response and exhibits a singularity $K^{-}_{\rm XC}\sim A(\omega)/q^{2}$ divergence at finite frequencies in the long-wavelength limit. The $1/q^{2}$ singularity in the exchange-correlation (XC) kernel is referred to as ultranonlocality in density functional theory (DFT). This concept describes how a finite change in the XC potential can be triggered by an infinitesimally small change in the density of the system Aschebrock et al. (2023). Specifically, the factor $A(\omega)$ is associated with the spin mass and the relaxation time as,
$A(\omega)=-\frac{m}{n\tau_{sd}}i\omega-\frac{m_{s}-m}{m}{n}{\omega^{2}}+\mathcal{O}(\omega^{3}),$ (2)
where $n$ is the density of electrons. This divergence is a direct manifestation of the non-conservation of spin currents and has significant consequences for the development of time-dependent density functional theory (TDDFT) Runge and Gross (1984); Elliott et al. (2009).
The ultranonlocality of the spin XC kernel poses severe challenges for the construction of approximate XC functional in TDDFT. Local or semi-local approximations, such as the adiabatic local spin density approximation Halilov et al. (1998), fail to capture the singular behavior of the kernel and the associated spin mass enhancement, leading to qualitatively incorrect predictions for spin transport. Advanced functionals that incorporate the ultranonlocal dependence of the spin currents and the spin Coulomb drag effect are therefore essential for the accurate description of spin dynamics in interacting systems.
Despite the fundamental importance of the spin mass, the spin diffusion, and their relation to the ultranonlocality of the spin XC kernel, direct numerical evidence of these phenomena has been lacking. Theoretical calculations have been limited to diagrammatic expansions D’Amico and Vignale (2000) and variational quantum Monte Carlo methods Qian and Vignale (2003); Ceperley and Alder (1980); Ortiz and Ballone (1994), which are restricted to the high-density regime and provide only approximate results. A quantitative understanding of the spin mass enhancement, the spin diffusion relaxation time, and their dependence on electron density has remained elusive, hindering the development of accurate XC functionals for spin transport.
In this paper, we present a systematic study of spin transport in interacting electron systems using the state-of-the-art Variational Diagrammatic Monte Carlo (VDMC) method Chen and Haule (2019); Haule and Chen (2022). Our calculations, spanning over a wide range of electron densities, momenta, and Matsubara frequencies, provide the first direct numerical evidence of the ultranonlocal behavior of the spin-resolved XC kernel, characterized by a $1/q^{2}$ divergence at finite frequencies. In addition, we demonstrate the absence of this divergence in the charge response, highlighting the unique nature of many-body effects in the spin channel. Furthermore, we extract precise values for the spin mass enhancement factor and the spin diffusion relaxation time, revealing their dependence on electron density and the suppression of the spin diffusion in the 3D uniform electrom gas (UEG) at low temperatures.
Our findings not only deepen the understanding of the interplay between XC effects and spin transport but also provide a quantitative benchmark for the development of advanced XC functionals that accurately capture the ultranonlocal behavior and the spin Coulomb drag effect. These functionals are essential for the reliable description of spin dynamics in interacting systems and the prediction of spin-dependent properties in real materials, which is essential for the further development of technologies in spintronics and quantum computing.
The implications of our work extend beyond the 3D UEG, paving the way for the exploration of spin transport in realistic systems, such as semiconductors, transition metals, and low-dimensional materials. The insights gained from our study can guide the design of novel spintronic devices that exploit the rich interplay between electron-electron interactions and spin dynamics, opening up new avenues for the manipulation and control of spin currents in quantum technologies.
The structure of our paper is as follows. In Sec. I, we introduce the theoretical background of spin transport in interacting electron systems, highlighting the importance of the spin Coulomb drag effect and its consequences, namely the spin mass enhancement and the spin diffusion phenomenon. Sec. II describes the VDMC method and its application to the calculation of the spin-spin response function in the 3D UEG. Our results are presented in two main sections: in Sec. III.1, we demonstrate the ultranonlocal behavior of the spin XC kernel and its absence in the charge response; in Sec. III.2 B, we extract the spin mass enhancement factor and the spin Coulomb drag relaxation time, discussing their dependence on electron density and the suppression of the dissipation of the spin currents at low temperatures. In Sec. IV, we summarize our conclusions and discuss the implications of our findings for the development of advanced XC functionals and the understanding of spin transport in realistic systems, as well as the potential impact on the design of spintronic devices.
@@ II Model & Method
Figure 1: Spin-spin response function $\chi_{S}$ at $q=0.3k_{\rm F}$ versus truncation order $N$ for $\theta=0.01$ and $r_{s}$ = 1. Panel (a) shows the static case, and Panel (b) shows the case with $\omega_{n}=0.503E_{\rm F}$. All $\lambda$ choices lead to the same extrapolated value, and the optimal $\lambda$ for the fastest convergence is about 2.0 for $\omega_{n}=0$ and 4.0 for $\omega_{n}=0.503E_{\rm F}$. Here $k_{\rm F}$ denotes the Fermi momentum, $E_{\rm F}$ denotes the Fermi energy and $N_{\rm F}$ denotes the density of state at the Fermi surface in the non-interacting system.
@@
Addressing the many-body problem of the UEG Hamiltonian poses significant challenges due to the divergences arising from the bare Coulomb interaction in the diagrammatic expansion Van Houcke et al. (2020). To overcome this issue, we employ the Variational Diagrammatic Monte Carlo (VDMC) method Chen and Haule (2019); Haule and Chen (2022); Prokof’ev and Svistunov (1998, 2008); Kozik et al. (2015); Rossi (2017); Prokof’ev and Svistunov (2007); Rossi et al. (2018); Van Houcke et al. (2012); Hou et al. (2024) , an advanced field-theoretic approach that offers controlled accuracy. The VDMC method transforms the problem into an equivalent and more appropriate form for the expansion, taking the emergent low-energy physics as the lowest order of the model. This transformation significantly improves the convergence of the expansion with increasing perturbation order.
Within the VDMC framework, the system’s action $S$ is decomposed into a reference action $S_{0}$ and a sequence of counterterms, serving as corrections. The Coulomb interaction inherent in the system is replaced by the Yukawa interaction $8\pi/(q^{2}+\lambda)$, with $\lambda$ serving as a variational screening parameter. This substitution allows for the representation of physical observables, such as electronic polarization, through a renormalized Feynman diagrammatic series Rossi et al. (2016); Peskin (2018), expanding in powers of the Yukawa term. The introduction of the “polarization” counterterm $\lambda/8\pi$ effectively cancels out large contributions arising from particle-hole fluctuations, expediting the convergence of the diagrammatic series. The parameter $\lambda$ is subject to iterative optimization to enhance convergence Kleinert (2009).
@@
High-order diagrams are efficiently evaluated through a Monte Carlo method employing importance sampling, with the sampling efficiency optimized using a computational graph representation of the diagrams Hou et al. (2024); Wang et al. (2021). The VDMC methodology has been successfully applied to explore various properties of the UEG, including the static and dynamic exchange-correlation kernel Chen and Haule (2019); Kukkonen and Chen (2021); LeBlanc et al. (2022), the effective mass Haule and Chen (2022), and the behavior of the electron gas under extreme conditions Hou et al. (2022). By optimizing the diagrammatic expansion, VDMC achieves reliable infinite-order results for any quantity without the need for a large truncation order $N$, significantly reducing computational costs while ensuring rapid and precise convergence for high-order calculations of physical observables.
In our investigation, we employ VDMC to evaluate the imaginary-time spin-spin correlation function $\chi_{S}(q,\tau)=\left<\mathcal{T}\hat{s}_{z}(q,\tau)\hat{s}_{z}(0)\right>$ in the thermodynamic limit. Subsequent Fourier transformation yields the correlation function in Matsubara frequency space. We then compare the dynamical spin correlation function with theoretical predictions given by Eq.(1) and Eq. (2) to probe the real-frequency dynamics of the spin-spin response function.
To validate the VDMC methodology, we calculate the spin-spin response function $\chi_{S}$ at $r_{s}=1,\theta=0.01,q=0.3k_{{\rm F}}$ up to the fifth diagrammatic order. Figure 1 illustrates rapid numerical convergence of $\chi_{S}$ in the vicinity of the optimal $\lambda$ for two characteristic Matsubara frequencies $\omega=0$ and $\omega=0.503E_{{\rm F}}$. This confirms VDMC’s capability for high-precision, high-order calculations.
Figure 2: Spin-spin response function $\chi_{S}$ as $\omega$ increases for different temperatures ($\theta=0.025,0.01$) with varying $q$ of the UEG system ($r_{s}=1.0$), indicating that the temperatures used in our simulations are effectively zero.
@@ III Results
We perform extensive VDMC simulations for 3D UEG systems with $r_{s}=0.5,1.0,2.0$ at temperatures $\theta=0.025,0.01$, measuring the spin-resolved response function $\chi_{S}$ for momenta $q$ within $0.5k_{\rm F}$ and Matsubara frequencies $\omega_{n}\leq E_{\rm F}$. Figure 2 shows the collapse of $\chi_{S}(i\omega_{n},\theta=0.025)$ and $\chi_{S}(i\omega_{n},\theta=0.01)$ onto the same curve for a given $q$, indicating that these temperatures are sufficiently low to converge to the zero-temperature limit. Consequently, we focus on simulations at $\theta=0.01$ to obtain a denser frequency grid for detailed study of the spin response characteristics.

### v4

# Matsubara-Frequency-Resolved Spin Exchange-Correlation Kernel for the Three-Dimensional Uniform Electron Gas

@@ Abstract
The spin Coulomb drag effect, arising from the exchange of momentum between electrons of opposite spins, plays a crucial role in the spin transport of interacting electron systems and can be characterized by the exchange-correlation (XC) kernel in the spin channel $K_{\rm XC}^{-}(q,\omega)$. Using the state-of-the-art Variational Diagrammatic Monte Carlo approach, we compute the Matsubara-frequency-resolved spin XC kernel $K_{\rm XC}^{-}(q,i\omega_{n})$ for the three-dimensional uniform electron gas at sufficiently low temperatures with high precision. In the long-wavelength limit, we identified a singular behavior of the form $A(i\omega_{n})/q^{2}$, confirming the theoretically predicted ultranonlocal behavior associated with spin Coulomb drag. Analysis of this structure in the low frequency region enables precise determination of two crucial parameters characterizing the spin Coulomb drag effect: the spin mass enhancement factor and spin diffusion relaxation time. We observe a significant trend of increasing enhancement of the spin mass factor with decreasing electron density, and provide clear evidence for the suppression of spin diffusion at low temperatures. These quantitative findings advance our understanding of Coulomb interaction effects on spin transport and provide essential parameters for time-dependent density functional theory and spintronics applications.
@@ I Introduction
Spin transport in interacting electron systems is a fundamental problem with far-reaching implications for spintronics and quantum technologies. The key challenge in describing spin transport lies in the complex interplay between electron-electron interactions and the non-conservation of spin currents. Unlike charge currents, which are conserved, spin currents can be dissipated through the exchange of momentum between electrons of opposite spins. As a result, the up-spin quasiparticle drags along some down-spin electrons in its “screening cloud”. This mechanism, known as the spin Coulomb drag (SCD) effect [1, 2], renders the spin dynamics fundamentally distinct from the charge dynamics.
The SCD effect has been extensively studied across various theoretical and experimental investigations in diverse systems, ranging from solid-state devices to cold atomic gases [3, 4, 5, 6, 7, 8, 9, 10]. A pioneering experimental confirmation was achieved by Weber et al. [5] in a GaAs quantum well, where measurements of spin trans-resistivity in quasi-two-dimensional electron gases exhibited quantitative agreement with theoretical predictions[6, 9, 10].
A crucial quantity for characterizing the SCD effect is the exchange-correlation (XC) kernel in the spin channel, defined as the difference between the inverse noninteracting susceptibility and the inverse spin susceptibility of the interacting electrons,
$K^{-}_{\mathrm{\rm XC}}(\mathbf{q},\omega)\equiv\chi^{-1}_{0}(\mathbf{q},\omega)-\chi^{-1}_{S}(\mathbf{q},\omega)$ (1)
with the superscript ‘--’ denoting its antisymmetric nature in the spin channel.
The XC kernel, parameterized from numerical calculations of the uniform electron gas (UEG) problem, plays a vital role in first-principles calculations of charge and spin properties in real materials. The static and uniform limit of the XC kernel provides a local spin-density approximation (LSDA) for density functional theory (DFT), essential for computing the thermodynamic magnetic properties of real materials [11, 12]. On the other hand, the frequency-resolved kernel is crucial for the time-dependent density functional theory (TDDFT) to incorporate electron correlations in ab-initio predictions of charge and spin dynamics [13]. Accurate parameterization of the XC kernel with the correct high-frequency and low-frequency limits is key to developing a TDDFT that captures qualitatively correct spin dynamics [14, 15].
In the long-wavelength limit, the spin XC kernel exhibits a singular behavior known as “ultranonlocality” [16, 17]. In this limit, both the interacting and non-interacting inverse spin susceptibilities exhibit a $1/q^{2}$-scaling. However, a striking difference emerges between the charge and spin XC kernel. In the charge sector, the conservation of charge currents ensures that the prefactors of this scaling are identical for both the interacting and non-interacting susceptibilities, leading to an exact cancellation in the charge XC kernel. In contrast, the absence of spin current conservation allows Coulomb interactions to renormalize the prefactor of the spin dynamical susceptibility. This renormalization prevents the cancellation that occurs in the charge sector, resulting in a singular behavior of the spin XC kernel,
$K_{XC}^{-}(\mathbf{q},\omega)=A(\omega)/q^{2}+O(1).$ (2)
Such singularity was first identified in the high-frequency limit through the third-moment sum rule [18, 19]. Subsequently, the low-frequency limit ($v_{F}q\ll\omega\ll E_{F}$ with $v_{F}$ and $E_{F}$ the Fermi velocity and the Fermi energy) of the XC kernel was conjectured based on the study of the SCD effect in spin current dynamics [1, 20, 21, 22]. The proposed low-frequency expansion is controlled by the spin mass and the relaxation time,
$A(\omega)=-\frac{m}{n\tau_{sd}}i\omega-\frac{m_{s}-m}{m}\frac{m}{n}{\omega^{2}}+\mathcal{O}(\omega^{3}),$ (3)
with $n$ the electron density and $m$ the electron mass, $\tau_{sd}$ is the spin relaxation time, and $m_{s}$ is the spin mass. The quantity $\tau_{sd}$ determines the lifetime of spin currents in the presence of electron-electron interactions and is essential for determining the spin diffusion length and overall efficiency of spin transport, which is proportional to the commonly studied SCD rate in previous work  [1, 6, 9, 10], while $m_{s}$ is a many-body parameter that quantifies the spin current carried by a single quasiparticle [20]. Both are two crucial many-body parameters for the SCD effect.
The ultranonlocality of the spin XC kernel poses significant challenges for accurate parameterization. Traditional methods, particularly the adiabatic local spin density approximation [23], fail to capture this ultranonlocal character, limiting their ability to describe spin-charge dynamics. To address this limitation, several studies [1, 6, 9] have proposed parameterizations based on insights gleaned from the random phase approximation (RPA) of current-current response functions. These response functions relate to the dynamic spin susceptibility through gauge symmetry. Notably, Reference [6] demonstrated that the SCD effect remains finite in the RPA of the current-current response function, even without taking into account XC corrections. This finding highlights the importance of the current-current response function as an alternative approach to studying the SCD effect. However, due to the lack of first-principles calculations of the spin XC kernel, these parameterizations rely on analytical ansatzes and approximations. While these approximations provide valuable insights, they may introduce systematic errors that are difficult to quantify.
To obtain a comprehensive understanding of the XC kernel, it is imperative to compute the frequency-resolved spin susceptibility for the UEG problem using first-principles numerical methods. While the static spin susceptibility in the UEG has been extensively studied with quantum Monte Carlo (QMC) methods [24, 25, 26, 27, 28, 29, 30, 31, 32, 33], the dynamic case remains far more challenging. Recent efforts have employed path-integral QMC to calculate the dynamic spin susceptibility in Matsubara frequency at high temperatures [34], but the accessible temperature range is insufficient to reveal the low-frequency structure of the spin XC kernel. This low-frequency behavior is crucial for understanding the SCD effect and developing reliable parameterizations for TDDFT calculations.
In this work, we develop a Variational Diagrammatic Monte Carlo (VDMC) method [35, 36] for calculating the dynamic spin susceptibility of the 3D UEG in the Matsubara frequency representation. By reaching sufficiently low temperatures, we precisely calculate the dynamical structure of the spin XC kernel $K_{\rm XC}^{-}(q,i\omega_{n})$ in both low and intermediate frequency regimes. Our numerical results across various electron densities confirm the theoretically predicted ultranonlocal singularity $A(i\omega_{n})/q^{2}$ in the long-wavelength limit. For comparison, we calculate the charge channel XC kernel $K_{\rm XC}^{+}(q,i\omega_{n})$, where this singularity is notably absent.
Furthermore, by performing least-squares fitting of our $K_{\rm XC}^{-}(q,i\omega_{n})$ data to an ansatz based on Eq. (2), we extract the low-frequency components of $A(i\omega_{n})$. The first-two order coefficients in the frequency expansion of $A(i\omega_{n})$ directly determine the spin mass enhancement and spin diffusion relaxation time, respectively, as shown in Eq. (3). Our calculations reveal a pronounced trend where the spin mass enhancement factor increases substantially as the electron density decreases, which are consistent with the previous results derived from the RPA methods but yield higher precisions [20]. Additionally, our results demonstrate vanishing inverse spin diffusion relaxation times, providing direct evidence for the suppression of spin diffusion processes at low temperatures as previous study has predicted [1]. Our study paves the way for the development of reliable ab initio methods for the spin dynamics in real materials.
The structure of our paper is as follows. Section II details the VDMC method for calculating the dynamic spin susceptibility in the 3D UEG. In section III.1, we demonstrate the ultranonlocal behavior of the spin XC kernel. In section III.2, we extract precise value of the spin mass enhancement factor and spin Coulomb drag relaxation time, discussing their dependence on electron density and temperature. Finally, Section IV summarizes our conclusions.
@@ II Model & Method
Figure 1: Spin susceptibility $\chi_{S}$ at $q=0.3k_{\rm F}$ versus truncation order $N$ for $\theta=0.01$ and $r_{s}$ = 1. Panel (a) shows the static case, and Panel (b) shows the case with $\omega_{n}=0.503E_{\rm F}$. All $\lambda$ choices lead to the same extrapolated value, and the optimal $\lambda$ for the fastest convergence is about 2.0 for $\omega_{n}=0$ and 4.0 for $\omega_{n}=0.503E_{\rm F}$. Here $k_{\rm F}$ denotes the Fermi momentum, $E_{\rm F}$ denotes the Fermi energy and $N_{\rm F}$ denotes the density of state at the Fermi surface in the non-interacting system.
@@
Addressing the many-body problem of the UEG Hamiltonian poses significant challenges due to the divergences arising from the bare Coulomb interaction in the diagrammatic expansion [37]. To overcome this issue, we employ the Variational Diagrammatic Monte Carlo (VDMC) method [35, 36, 38, 39, 40, 41, 42, 43, 44, 45] , an advanced field-theoretic approach that offers controlled accuracy. The VDMC method transforms the problem into an equivalent and more appropriate form for the expansion, taking the emergent low-energy physics as the lowest order of the model. This transformation significantly improves the convergence of the expansion with increasing perturbation order.
Within the VDMC framework, the system’s action $S$ is decomposed into a reference action $S_{0}$ and a sequence of counterterms, serving as corrections. The Coulomb interaction inherent in the system is replaced by the Yukawa interaction $8\pi/(q^{2}+\lambda)$, with $\lambda$ serving as a variational screening parameter. This substitution allows for the representation of physical observables, such as electronic polarization, through a renormalized Feynman diagrammatic series [46, 47] , expanding in powers of the Yukawa term. The introduction of the “polarization” counterterm $\lambda/8\pi$ effectively cancels out large contributions arising from particle-hole fluctuations, expediting the convergence of the diagrammatic series. The parameter $\lambda$ is subject to iterative optimization to enhance convergence [48].
@@
High-order diagrams are efficiently evaluated through a Monte Carlo method employing importance sampling, with the sampling efficiency optimized using a computational graph representation of the diagrams [45, 49]. The VDMC methodology has been successfully applied to explore various properties of the UEG, including the static and dynamic exchange-correlation kernel [35, 32, 50], the effective mass [36], and the behavior of the electron gas under extreme conditions [51]. By optimizing the diagrammatic expansion, VDMC achieves reliable infinite-order results for any quantity without the need for a large truncation order $N$, significantly reducing computational costs while ensuring rapid and precise convergence for high-order calculations of physical observables.
In our investigation, we employ VDMC to evaluate the imaginary-time spin-spin correlation function $\chi_{S}(q,\tau)=\left<\mathcal{T}\hat{s}_{z}(q,\tau)\hat{s}_{z}(0)\right>$ in the thermodynamic limit. Subsequent Fourier transformation yields the correlation function in Matsubara frequency space. We then compare the dynamical spin correlation function with theoretical predictions given by Eqs. (1) - (3) to probe the real-frequency dynamics of the spin susceptibility.
To validate the VDMC methodology, we calculate the spin susceptibility $\chi_{S}$ at $r_{s}=1,\theta=0.01,q=0.3k_{{\rm F}}$ up to the fifth diagrammatic order. Figure 1 illustrates rapid numerical convergence of $\chi_{S}$ in the vicinity of the optimal $\lambda$ for two characteristic Matsubara frequencies $\omega_{n}=0$ and $\omega_{n}=0.503E_{{\rm F}}$. This confirms VDMC’s capability for high-precision, high-order calculations.
Figure 2: Dynamic spin susceptibility $\chi_{S}$ as the Matsubara frequency $\omega_{n}$ increases for different temperatures ($\theta=0.025,0.01$) with varying $q$ of the UEG system ($r_{s}=1.0$), indicating that the temperatures used in our simulations are effectively zero.
@@ III Results
We perform extensive VDMC simulations for 3D UEG systems with $r_{s}=0.5,1.0,2.0$ at temperatures $\theta=0.025,0.01$, measuring the spin susceptibility $\chi_{S}$ for momenta $q$ within $0.5k_{\rm F}$ and Matsubara frequencies $\omega_{n}\leq E_{\rm F}$. Figure 2 shows the collapse of $\chi_{S}(i\omega_{n},\theta=0.025)$ and $\chi_{S}(i\omega_{n},\theta=0.01)$ onto the same curve for a given $q$, indicating that these temperatures are sufficiently low to converge to the zero-temperature limit. Consequently, we focus on simulations at $\theta=0.01$ to obtain a denser frequency grid for detailed study of the spin response characteristics.